OpenStudy (anonymous):

The volume of a box is 2176 cubic feet. The width of the box is 8 feet and it's height is 1 foot more than it's length. Find the height of the box. can anyone solve. Please show work step by step so I can understand.

OpenStudy (anonymous):

formula for volume = lwh given that h= l+1 it will be --- 2176 = 8(L)(L+1)

OpenStudy (anonymous):

can you do it now?

OpenStudy (anonymous):

what is the length?

OpenStudy (anonymous):

you have to find length first, using quadratic formula, after getting length(L) add it with 1 as stated that height is 1 foot more than it's length. .

OpenStudy (anonymous):

okay. what is a b and c. I know the formula. but not how to apply it to the problem

OpenStudy (anonymous):

the equation will be. l^2 + l - 272 = 0 a=1 b= 1 c= -272

OpenStudy (anonymous):

ok. i think i understand. thanks

OpenStudy (anonymous):

glad to help. .

Directrix (directrix):

2176 = 8 L ( 1 + L) 2176 = 8L + 8 L^2 8 L^2 + 8L - 2176 - 0 8 (L^2 + L - 272) = 0 (L +17)(L -16) = 0 L = - 17 or L = 16 L = 16 h = L + 1 = 17